OpenStudy (anonymous):

what is the percent of dercrease from 80 to 52

OpenStudy (anonymous):

52/80=0.65

OpenStudy (anonymous):

so 100-65%=35

OpenStudy (beccaboo333):

Please do not just give the answer, have them work through an equation.

OpenStudy (beccaboo333):

@lucyv7129

OpenStudy (anonymous):

ok, sorry, but that's right, right?

OpenStudy (texaschic101):

percent decrease formula : (older - newer) / older....x 100 (80 - 52) / 80 = 28/80 = 0.35 0.35 x 100 = 35% lucyv....you got the correct answer...I just do it a different way :)